OpenStudy (anonymous):

Find the exact value of each trigonometric function. Some may be undefined. csc 5π/3

OpenStudy (anonymous):

Here my work. r/y= 1/√3/2= √3/2

OpenStudy (anonymous):

it looks wrong

OpenStudy (anonymous):

yeah it is wrong what you need is \[\sin(\frac{5\pi}{3})\] and then take the reciprocal

OpenStudy (anonymous):

you are close, but \(\sin(\frac{5\pi}{3})\) is negative

OpenStudy (anonymous):

you got it now?

OpenStudy (anonymous):

oh also another mistake. the reciprocal of \(-\frac{\sqrt{3}}{2}\) is \(-\frac{2}{\sqrt{3}}\)

OpenStudy (anonymous):

seems like in your first post you forgot to flip it